OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

It would be the middle selection, b^2 - 4ac because when that is an integer squared, then you can easily factor. As for getting the value of b^2 - 4ac in this case, just substitute from your equation which is in the form ax^2 + bx + c = 0. So, a=2, b=7, and c=3.

OpenStudy (anonymous):

Oh ok, thx so much! :)

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Thx again! :D P.S. is x a?

OpenStudy (anonymous):

Not quite sure what the question is here. Is it : is x = a?

OpenStudy (anonymous):

Yep! That is what I mean. :)

OpenStudy (anonymous):

np. What you are essentially asking is: How do we factor the given equation. That will determine x. So, 2x2 + 7x + 3 = 0 -> (2x + 1)(x + 3) = 0 So, x will equal -1/2 or -3. That's for x. As for what b^2 - 4ac, that is called the discriminant because when that is positive or 0, the equation will have real solutions. If the discriminant is an integer squared, like here it is 25 which is 5^2, then we know we can factor. Does this help?

OpenStudy (anonymous):

Yep, thx!!! :D

OpenStudy (anonymous):

Ok, that's all I need! :D

OpenStudy (anonymous):

You're quite welcome!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I don't think I need anything else1 I totally understand this lesson now, you have been very helpful! :D